OpenStudy (anonymous):

Find the x and y intercepts. I just need to make sure this is right please! 4x-7y=84 x intercept should be (21,0) and the y intercept should be (0,12)

OpenStudy (anonymous):

Correct!!

OpenStudy (anonymous):

y=-12

OpenStudy (anonymous):

coz for calculating y intercept put x=0 in the above equation 4x-7y=84 -7y=84 so y=-84/7 y=-12